OpenStudy (soccergirlforever1):

three times the square of a non-zero number is equal to fifteen times the number.What is the number?

OpenStudy (texaschic101):

3x^2 = 15x 3x^2 - 15x = 0 3x (x - 5) = 0 3x = 0 x = 0....disregard this...it can be eliminated x - 5 = 0 x = 5...this is ur number